Water is falling on the blades of a turbine at a rate 6000Kg/min. The height of the fall is100m. What is the power gained by the turbine?
A. 10KW B. 6KW C. 100KW D. 600KW

Grade:

1 Answers

Ishan
7 years ago
M = 6 x 10^ 3h = 100mg = 10m/spower = rate of doing workt = 60spower = mgh/tpower= 6 x 10^3 x 10 x 100/ 60power = 10^5 = 100000 W = 100kW.So, the power gained by the turbine is 100kW.Hence, option c is correct.Thank you!